# TabiBERT:A Large-Scale ModernBERT Foundation Model and Unified Benchmarking Framework for Turkish

### v1

@@ Abstract
The introduction of BERT established encoder-only Transformer models as a foundational paradigm in natural language processing, motivating extensive research on architectural design and pre-training strategies. Subsequent work addressed limitations of early encoder models through improvements in computational efficiency, training stability, context length, and positional encoding. ModernBERT consolidated these improvements by integrating rotary positional embeddings, FlashAttention-based attention implementations, and refined normalization schemes to support longer contexts and more stable training. Unfortunately, the Turkish NLP landscape lacks a monolingual encoder trained from scratch that systematically incorporates these modern architectural advances.
This work introduces TabiBERT, a monolingual Turkish encoder based on the ModernBERT architecture that is trained from scratch on a large, carefully curated corpus. TabiBERT is pre-trained on one trillion tokens sampled from an 84.88B token multi-domain corpus comprising web text (73%), scientific publications (20%), source code (6%), and mathematical content (0.3%). The model supports a context length of 8,192 tokens, which is 16x of the original BERT models. It achieves an inference speedup of up to 2.65 times while reducing GPU memory consumption – allowing for larger batch sizes during training and inference. To support rigorous and reproducible evaluation, we introduce the TabiBench benchmark with 28 datasets across eight task categories with standardized splits and evaluation protocols. Performance is summarized using a GLUE-style macro-average score. On TabiBench, TabiBERT attains a score of 77.58, outperforming BERTurk by 1.62 points and establishing new state-of-the-art results on five of eight task categories, with particularly strong gains on question answering (+9.55 points), code retrieval (+2.41 points), and document retrieval (+0.60 points). When compared with task-specific prior best results, including specialized models such as TurkishBERTweet, TabiBERT achieves an average improvement of 1.47 points, indicating robust cross-domain generalization. We release model weights, training configurations, and evaluation code to provide a transparent and reproducible foundation for future Turkish encoder research.
@@ 1 Introduction
The emergence of pre-trained transformer encoder models has become central to contemporary natural language processing, providing general-purpose representations that support a wide range of downstream tasks. The architectural framework introduced by Bidirectional Encoder Representations from Transformers (BERT) (Devlin2019BERT) has been a foundational component of natural language understanding, enabling the induction of deeply contextualized linguistic features through self-supervised learning on large-scale corpora. BERT was first trained on large English text corpora and achieved strong baseline performance on a wide range of natural language processing tasks, serving as the basis for many later pre-trained language models.
@@
The success of BERT motivated, and thus was followed by extensive research, leading to resource-efficient variants (Liu2019RoBERTa; Sanh2019DistilBERT; Jiang2020ConvBERT; Lan2020ALBERT; Clark2020ELECTRA), domain-specialized models (Beltagy2019SciBERT; Huang2019ClinicalBERT; Lee2020BioBERT; Chalkidis2020LegalBERT), and multilingual architectures designed to support cross-lingual transfer (Pires2019MultilingualBERT; Vulic2020ContinualLearning; Conneau2020XLM-R). Initial refinements focused on pre-training procedures that improved performance without changing the core architecture. For example, RoBERTa (Liu2019RoBERTa) achieved improved performance via larger and more diverse pre-training corpora, dynamic masking strategies, optimized training regimes, and by eliminating the next sentence prediction objective without altering the original transformer architecture. Several works concurrently focused on efficiency through knowledge distillation (Sanh2019DistilBERT), parameter sharing strategies (Lan2020ALBERT), and alternative pre-training objectives (Clark2020ELECTRA), enabling computational gains while maintaining or improving performance.
@@
Subsequent architectural innovations sought to address BERT’s quadratic complexity, which stems from the self-attention mechanism and limits sequence length and computational efficiency. Beltagy2020Longformer introduced sparse attention to enable longer sequences, while IO-aware implementations significantly reduced memory overhead (Dao2022FlashAttention). Positional encoding improvements such as Rotary Position Embeddings (RoPE) (Su2024Roformer) resulted in improved length extrapolation compared to BERT’s original absolute positional embeddings. Normalization strategies were introduced to improve training stability for deeper architectures (Xiong2020LayerNorm; Zhang2019RootMeanSquare). These accumulated improvements were synthesized into ModernBERT, an updated encoder-only architecture that incorporates rotary positional embeddings, FlashAttention technology, alternating attention patterns, and refined pre-training procedures (Warner2024ModernBERT). Its architecture supports context windows up to $8,192$ tokens, sixteen times BERT’s original context length of 512, with improved computational efficiency that addresses limitations stemming from fine-tuning instability and restricted sequence length (Xiong2020LayerNorm; Liu2019RoBERTa).
@@
While these architectural advances have been widely adopted for high-resource languages, Turkish NLP remains primarily centered on earlier architectural paradigms. BERTurk (Schweter2020BERTurk) laid the foundation for transformer-based Turkish NLP as the first monolingual BERT model pre-trained from scratch and publicly released. Its widespread adoption across diverse downstream tasks has made it the de facto standard for Turkish language understanding. Following, subsequent models such as ConvBERTurk and various ELECTRA adaptations have emerged on community platforms, although they frequently prioritized immediate availability over technical documentation. Critical specifications regarding corpus composition, preprocessing pipelines, and systematic evaluation protocols are often underspecified, which impedes reproducibility and limits the ability of the research community to build systematically on prior work.
@@
Recent large-scale multilingual efforts have incorporated Turkish alongside other languages. Multilingual BERT (mBERT) (Devlin2019BERT) and XLM-R (Conneau2020XLM-R) share parameters and vocabulary across typologically diverse languages, which dilutes representational capacity for individual languages (Conneau2020XLM-R; Vulic2020ContinualLearning). For morphologically rich languages, multilingual tokenizers produce inefficient segmentations, leading to increased token fertility, longer effective sequences, and reduced vocabulary utilization (Rust2021Tokenizer; Hofmann2022Tokenization).
@@
Model # Params (M) Data Size (GB) Context Length Code & Math Flash Attn. TabiBench Score
@@
We address these gaps by introducing TabiBERT, a monolingual Turkish encoder based on the ModernBERT architecture, and TabiBench, a unified evaluation benchmark with standardized protocols. TabiBERT is a monolingual Turkish encoder based on the ModernBERT architecture, trained from scratch on a large, carefully curated corpus. Across TabiBench, TabiBERT achieves an average score of 77.58 compared to the next best model BERTurk’s 75.96, representing an improvement of 1.62 absolute points (on a 0-100 scale, following GLUE-style macro-averaging) and establishing state-of-the-art results among Turkish-specific encoders on five of eight task categories. Performance gains are particularly substantial on retrieval tasks, with improvements of 9.55 points on question answering, 2.41 points on code retrieval, as well as 0.60 points on document retrieval. When compared against the previous best result for each task individually, TabiBERT achieves an average improvement of 1.47 points, demonstrating strong cross-domain generalization with a single general-purpose architecture. Beyond accuracy improvements, TabiBERT provides 2.65× faster inference and supports a context length of 8,192 tokens, sixteen times that of prior Turkish BERT models, while reducing GPU memory consumption to enable larger batch sizes during training and inference.
More recently, mmBERT (Marone2025mmBERT), a 307307M-parameter multilingual model with a vocabulary of $256,000$ tokens trained from scratch at large scale, demonstrates strong performance on Turkish tasks, achieving 79.26 on TabiBench (see Appendix A). However, this performance comes at the cost of significantly increased computational requirements and the need for large-scale multilingual training data. Moreover, mmBERT’s multilingual tokenizer produces substantially higher token fertility for Turkish texts, +41% compared to TabiBERT’s monolingual tokenizer, resulting in longer effective sequences and reduced computational efficiency for Turkish-specific applications (detailed in Section 5). We include mmBERT as a reference point representing a different model class with substantially different resource requirements and design objectives, while our work focuses on developing an efficient, monolingual encoder that incorporates modern architectural advances at a more modest scale suitable for research and practical deployment.
@@
  • TabiBERT (TabiBERT-HF), a state-of-the-art Turkish encoder that bridges the architectural gap in the Turkish NLP ecosystem by adapting the ModernBERT architecture. The model was pre-trained for a total of 1 trillion tokens on a multi-domain Turkish corpus comprising 84.8884.88B tokens, including web text (73%), scientific publications (20%), source code (6%), and mathematical content (0.3%). TabiBERT features an $8,192$-token context length, representing a 15× increase over previous Turkish BERT-style models, and achieves up to 2.65× inference speedup. The model achieves state-of-the-art results among Turkish-specific encoders on five of eight benchmark task categories, with an average improvement of 1.62 absolute points over BERTurk and particularly strong gains on question answering (9.55 points) and retrieval tasks (2.41 points on code retrieval, 0.60 points on document retrieval), while significantly reducing memory overhead.
  • TabiBench (TabiBench), a unified and comprehensive evaluation framework designed to standardize Turkish NLU research by providing a reproducible pipeline including standardized data splits, preprocessing scripts, and evaluation configurations. TabiBench comprises 28 datasets across 8 task categories, including specialized domains such as code retrieval and academic text understanding that have not been systematically evaluated in prior Turkish NLP benchmarks.
  • The open-source release of the TabiBERT model weights and the TabiBench evaluation codebase (TabiBERT) to provide a transparent foundation for future research in Turkish language models and downstream applications. Detailed corpus composition statistics and preprocessing methodology are provided in Section 3.
The remainder of the paper is structured as follows: Section 2 reviews the relevant literature on encoder architectures and Turkish language models; Section 3 presents the corpus and outlines the data preparation procedures; Section 4 details the pre-training methodology and architectural choices; Section 5 reports a comprehensive evaluation of TabiBERT across the TabiBench benchmark; and Section 6 summarizes the findings, discusses limitations, and identifies directions for future research.
@@ 2 Related Work
The introduction of Bidirectional Encoder Representations from Transformers (BERT) (Devlin2019BERT) was a significant advancement in encoder-only transformer models. These models were trained on large English corpora and established strong performance baselines across diverse natural language processing tasks. They provided a pre-trained model that could be fine-tuned for various downstream tasks.
@@ 2.1 Encoder-only Transformer Models & Architectural Evolution
Following BERT, encoder-only transformer models were improved by various architectural innovations or training techniques, including changes to pre-training objectives, optimization strategies, and large-scale data and model scaling, leading to consistent performance gains across a wide range of downstream NLP tasks. RoBERTa (Liu2019RoBERTa) was trained with the same masked language modeling objective but with a larger training corpus and bigger model sizes, achieving state-of-the-art performance on various NLP tasks. Following, ALBERT (Lan2020ALBERT) introduced a number of mechanisms such as parameter sharing and a factorized embedding parameterization, resulting in a significant reduction in model size and memory consumption. This enabled training deeper, wider encoder models while still performing competitively on downstream benchmarks. ELECTRA (Clark2020ELECTRA) proposed a more sample-efficient pre-training approach based on replaced token detection, where a discriminator model learns to distinguish real input tokens from plausible alternatives generated by a small generator network, resulting in strong performance with substantially reduced pre-training compute. DeBERTa (he2021debertav1) introduced disentangled attention to decouple content and positional representations, improving modeling of relative position information in encoder-only transformers. The success of DeBERTa was followed by DeBERTaV2 (microsoft_debertav2_repo), which refined this approach by using improved mask decoding and pre-training strategies. Moreover, DeBERTaV3 (He2023DeBERTaV3) further enhanced training efficiency by integrating a replaced-token detection objective inspired by ELECTRA.
@@
Subsequent encoder-only transformer works after DeBERTaV3 focused on improving pre-training efficiency by revisiting training objectives and modernizing the BERT architecture rather than attempting to introduce novel attention mechanisms. MosaicBERT (Portes2023Mosaicbert) worked on engineering a training recipe by combining architectural and systems optimizations such as Flash Attention (Dao2022FlashAttention), ALiBi positional biases (Press2021Train), Gated Linear Units (Shazeer2020Glu), dynamic unpadding (Zeng2022Boosting), and low-precision normalization. This engineering effort demonstrated that BERT pre-training can run significantly faster and more cost-effectively, while maintaining competitive downstream performance. In parallel, task-oriented pre-training strategies were explored, exemplified by RetroMAE (Xiao2022RetroMAE), which redesigned masked autoencoding to better support dense retrieval tasks using a BERT-style encoder. FlexiBERT (Tuli2023FlexiBERT) used neural architecture search to identify performant encoder configurations under different computational constraints, aiming to investigate architectural flexibility and efficiency trade-offs. Renewed interest in encoder-only models has led to modernized successors to BERT. ModernBERT (Warner2024ModernBERT) revisited the original BERT model, extending context length and updating architectural elements and design choices. NeoBERT (Breton2025NeoBERT) and OptiBERT (Dervishi2025TrainingOptiBERT) emphasized compute-optimal scaling and practical deployment. Together, these studies show a trend toward improving the efficiency and practicality of BERT-style encoder models, while maintaining strong performance on language understanding tasks.
@@ 2.2 Monolingual Turkish BERT Models
The success of transformer-encoder models motivated a series of BERT models known as BERTurk to be trained on Turkish data (Schweter2020BERTurk). Several variations of BERT, such as DistilBERT (Sanh2019DistilBERT), ConvBERTurk (Schweter2020BERTurkV2), and electra (Clark2020ELECTRA) were also trained on Turkish data by the same authors (Schweter2020BERTurk). These models were generally trained on a 3535GB corpus consisting of 4.44.4B tokens drawn from the Turkish oscar corpus (Abadji2022OSCAR), a Wikipedia dump [^1] An unspecified Wikipedia dump, presumed to date from approximately 2020., and various OPUS corpora (Tiedemann2012OPUS). Some models, like ConvBERTurk and electra, were also trained on the Turkish portion of the mC4 corpus (xue2021mt5massivelymultilingualpretrained; Schweter2020BERTurk) which is a cleaned version of the public web crawl data of Common Crawl. These models have been evaluated on various downstream tasks (such as part-of-speech tagging, named entity recognition, and question answering) where they generally outperform their multilingual counterparts (Schweter2020BERTurk; wu2019betobentzbecassurprising; Conneau2020XLM-R). Despite their popularity, these models were limited in transparency as they were not products of academic publications.
Following, two Turkish BERT models were introduced with academic publications. YTU-Cosmos-BERT (Kesgin2023YTU) utilized an uncased Turkish tokenizer and was trained on a corpus of an unknown number of tokens reported to be around 7575GB in size, that consists of text collected from the web, Turkish Wikipedia and novels. TurkishBERTweet (Najafi2023TurkishBERTweet) was tailored for social media. It also employed an uncased tokenizer, specifically designed for social media texts, by the addition of special tokens that can encode special information in tweets like hashtags, mentions, and URLs. It was pre-trained with a corpus of around 110110GB in uncompressed form, containing about 900 million tweets of 10 tokens on average, which totals to a sum of 99B tokens when extrapolated. Recently, a RoBERTa-based model for Turkish was published (ScheibleSchmitt2025SindBERT), which has been an improvement over the previous RoBERTa-based models for Turkish (TasRoberTURK2024) as it was trained on a larger corpus, using a larger batch size. However, it did not achieve a significant improvement over the previous models.

### v3

# TabiBERT:A Large-Scale ModernBERT Foundation Model and A Unified Benchmark for Turkish

@@ Abstract
The introduction of BERT established encoder-only Transformer models as a foundational paradigm in natural language processing, motivating extensive research on architectural design and pretraining strategies. Subsequent work addressed limitations of early encoder models through improvements in computational efficiency, training stability, context length, and positional encoding. ModernBERT consolidated these improvements by integrating rotary positional embeddings, FlashAttention-based attention implementations, and refined normalization schemes to support longer contexts and more stable training. Unfortunately, the Turkish NLP landscape lacks a monolingual encoder trained from scratch that systematically incorporates these modern architectural advances.
This work introduces TabiBERT, a monolingual Turkish encoder based on the ModernBERT architecture that is trained from scratch on a large, carefully curated corpus. TabiBERT is pretrained on one trillion tokens sampled from an 84.88B token multi-domain corpus comprising web text (73%), scientific publications (20%), source code (6%), and mathematical content (0.3%). The model supports a context length of 8,192 tokens, which is 16x of the original BERT models. It achieves an inference speedup of up to 2.65x while reducing GPU memory consumption, allowing for larger batch sizes during training and inference. To support rigorous and reproducible evaluation, we introduce the TabiBench benchmark with 28 datasets across eight task categories with standardized splits and evaluation protocols. Performance is summarized using a GLUE-style macro-average score. On TabiBench, TabiBERT attains a score of 77.58, outperforming BERTurk by 1.62 points and establishing new state-of-the-art results on five of eight task categories, with particularly strong gains on question answering (+9.55 points), code retrieval (+2.41 points), and academic understanding (+0.66 points). When compared with task-specific prior best results, including specialized models such as TurkishBERTweet, TabiBERT achieves an average improvement of 1.47 points, indicating robust cross-domain generalization. We release model weights, training configurations, and evaluation code to provide a transparent and reproducible foundation for future Turkish encoder research.
@@ 1 Introduction
The emergence of pretrained transformer encoder models has become central to contemporary natural language processing, providing general-purpose representations that support a wide range of downstream tasks. The architectural framework introduced by Bidirectional Encoder Representations from Transformers (BERT) (Devlin2019BERT) has been a foundational component of natural language understanding, enabling the induction of deeply contextualized linguistic features through self-supervised learning on large-scale corpora. BERT was first trained on large English text corpora and achieved strong baseline performance on a wide range of natural language processing tasks, serving as the basis for many later pretrained language models.
@@
The success of BERT motivated, and thus was followed by extensive research, leading to resource-efficient variants (Liu2019RoBERTa; Sanh2019DistilBERT; Jiang2020ConvBERT; Lan2020ALBERT; Clark2020ELECTRA), domain-specialized models (Beltagy2019SciBERT; Huang2019ClinicalBERT; Lee2020BioBERT; Chalkidis2020LegalBERT), and multilingual architectures designed to support cross-lingual transfer (Pires2019MultilingualBERT; Vulic2020ContinualLearning; Conneau2020XLM-R). Initial refinements focused on pretraining procedures that improved performance without changing the core architecture. For example, RoBERTa (Liu2019RoBERTa) achieved improved performance via larger and more diverse pretraining corpora, dynamic masking strategies, optimized training regimes, and by eliminating the next sentence prediction objective without altering the original transformer architecture. Several works concurrently focused on efficiency through knowledge distillation (Sanh2019DistilBERT), parameter sharing strategies (Lan2020ALBERT), and alternative pretraining objectives (Clark2020ELECTRA), enabling computational gains while maintaining or improving performance.
@@
Subsequent architectural innovations sought to address BERT’s quadratic complexity, which stems from the self-attention mechanism and limits sequence length and computational efficiency. Beltagy2020Longformer introduced sparse attention to enable longer sequences, while IO-aware implementations significantly reduced memory overhead (Dao2022FlashAttention). Positional encoding improvements, such as Rotary Position Embeddings (RoPE) (Su2024Roformer) resulted in improved length extrapolation compared to BERT’s original absolute positional embeddings. Normalization strategies were introduced to improve training stability for deeper architectures (Xiong2020LayerNorm; Zhang2019RootMeanSquare). These accumulated improvements were synthesized into ModernBERT, an updated encoder-only architecture that incorporates rotary positional embeddings, FlashAttention technology, alternating attention patterns, and refined pretraining procedures (Warner2024ModernBERT). Its architecture supports context windows up to $8,192$ tokens, sixteen times BERT’s original context length of 512, with improved computational efficiency that addresses limitations stemming from fine-tuning instability and restricted sequence length (Xiong2020LayerNorm; Liu2019RoBERTa).
@@
While these architectural advances have been widely adopted for high-resource languages, Turkish NLP remains primarily centered on earlier architectural paradigms. BERTurk (Schweter2020BERTurk) laid the foundation for transformer-based Turkish NLP as the first monolingual BERT model pretrained from scratch and publicly released. Its widespread adoption across diverse downstream tasks has made it the de facto standard for Turkish language understanding. Following, subsequent models such as ConvBERTurk and various ELECTRA adaptations have emerged on community platforms, although they frequently prioritized immediate availability over technical documentation. Critical specifications regarding corpus composition, preprocessing pipelines, and systematic evaluation protocols are often underspecified, which impedes reproducibility and limits the ability of the research community to build systematically on prior work.
@@
Recent large-scale multilingual efforts have incorporated Turkish alongside other languages. Multilingual BERT (mBERT) (Devlin2019BERT) and XLM-R (Conneau2020XLM-R) share parameters and vocabulary across typologically diverse languages, which dilutes representational capacity for individual languages (Conneau2020XLM-R; Vulic2020ContinualLearning). For morphologically rich languages, multilingual tokenizers produce inefficient segmentations, leading to increased token fertility, longer effective sequences, and reduced vocabulary utilization (Rust2021HowGood; Hofmann2022Tokenization).
@@
Model # Params (M) Data Size (GB) Context Length Code & Math Flash Attention TabiBench Score
@@
We address these gaps by introducing TabiBERT, a monolingual Turkish encoder, based on the ModernBERT architecture, and TabiBench, a unified evaluation benchmark with standardized protocols. TabiBERT is a monolingual Turkish encoder based on the ModernBERT architecture, trained from scratch on a large, carefully curated corpus. Across TabiBench, TabiBERT achieves an average score of 77.58 compared to the next best model BERTurk’s 75.96, representing an improvement of 1.62 absolute points (on a 0-100 scale, following GLUE-style macro-averaging) and establishing state-of-the-art results among Turkish-specific encoders on five out of eight task categories. TabiBERT’s technical specifications and results compared to prior Turkish BERT models are summarized in Table 1. Performance gains are particularly substantial on four tasks, with improvements of 9.55 points on question answering, 2.41 points on code retrieval, 0.66 points on academic understanding, as well as 0.60 points on information retrieval. When compared against the previous best result for each task individually, TabiBERT achieves an average improvement of 1.47 points, demonstrating strong cross-domain generalization with a single general-purpose architecture. Beyond accuracy improvements, TabiBERT provides 2.65× faster inference and supports a context length of 8,192 tokens, sixteen times that of prior Turkish BERT models, while reducing GPU memory consumption to enable larger batch sizes during training and inference.
For completeness, we also evaluate against mmBERT (Marone2025mmBERT), a recent large-scale multilingual encoder with 307307M parameters and a $256,000$-token vocabulary that achieves 79.26 on TabiBench. mmBERT represents a different model class, trained on multilingual data at substantially larger scale with correspondingly higher computational requirements. We include it as a reference point to situate TabiBERT within the broader landscape of models applicable to Turkish. Notably, mmBERT’s multilingual tokenizer has 41 per cent higher token fertility for Turkish texts compared to TabiBERT’s monolingual tokenizer, resulting in longer effective sequences and reduced computational efficiency for Turkish-specific applications (detailed in Section 5). Our work focuses on developing an efficient, monolingual encoder that incorporates modern architectural advances at a scale suitable for widespread research and practical deployment.
@@
  • TabiBERT (TabiBERT-HF), a state-of-the-art monolingual Turkish encoder based on the ModernBERT architecture. The model was pretrained for 1T tokens on a corpus spanning diverse domains and achieves state-of-the-art results among Turkish-specific encoders on five of eight benchmark categories. TabiBERT provides substantial efficiency improvements, including 2.65× faster inference, 16× longer context length (8,192 tokens), and reduced memory overhead compared to prior Turkish BERT models.
  • TabiBench (TabiBench), a unified evaluation framework comprising 28 datasets across 8 task categories with standardized protocols, data splits, and preprocessing pipelines. The benchmark includes specialized domains such as code retrieval and academic text understanding that extend beyond prior Turkish NLP evaluation frameworks.
  • Open-source release of model weights and evaluation code (TabiBERT) to provide a reproducible foundation for Turkish NLP research.
The remainder of this paper is organized as follows: Section 2 reviews prior work on encoder architectures, Turkish language models and evaluation benchmarks, Section 3 describes the TabiBERT model and the pretraining procedure, Section 4 specifies the TabiBench evaluation framework, Section 5 reports the evaluation results of TabiBERT on the TabiBench benchmark, Section 6 discusses the results, Section 7 presents some limitations of our work, and Section 8 concludes the paper and outlines directions for future work.
@@ 2 Related Work
The introduction of Bidirectional Encoder Representations from Transformers (BERT) (Devlin2019BERT) was a significant advancement in encoder-only transformer models. These models were trained on large English corpora and established strong performance baselines across diverse natural language processing tasks. They provided a pretrained model that could be fine-tuned for various downstream tasks.
@@ 2.1 Encoder-only Transformer Models & Architectural Evolution
Following BERT, encoder-only transformer models were improved by various architectural innovations or training techniques, including changes to pretraining objectives, optimization strategies, and large-scale data and model scaling, leading to consistent performance gains across a wide range of downstream NLP tasks. RoBERTa (Liu2019RoBERTa) was trained with the same masked language modeling objective but with a larger training corpus and bigger model sizes, achieving state-of-the-art performance on various NLP tasks. Following, ALBERT (Lan2020ALBERT) introduced a number of mechanisms such as parameter sharing and a factorized embedding parameterization, resulting in a significant reduction in model size and memory consumption. This enabled training deeper, wider encoder models while still performing competitively on downstream benchmarks. ELECTRA (Clark2020ELECTRA) proposed a more sample-efficient pretraining approach based on replaced token detection, where a discriminator model learns to distinguish real input tokens from plausible alternatives generated by a small generator network, resulting in strong performance with substantially reduced pretraining compute. DeBERTa (he2021debertav1) introduced disentangled attention to decouple content and positional representations, improving modeling of relative position information in encoder-only transformers. The success of DeBERTa was followed by DeBERTaV2 (MicrosoftDeBERTaV2Repo), which refined this approach by using improved mask decoding and pretraining strategies. Moreover, DeBERTaV3 (He2023DeBERTaV3) further enhanced training efficiency by integrating a replaced-token detection objective inspired by ELECTRA.
@@
Subsequent encoder-only transformer works after DeBERTaV3 focused on improving pretraining efficiency by revisiting training objectives and modernizing the BERT architecture rather than attempting to introduce novel attention mechanisms. MosaicBERT (Portes2023Mosaicbert) worked on engineering a training recipe by combining architectural and systems optimizations such as Flash Attention (Dao2022FlashAttention), ALiBi positional biases (Press2021Train), Gated Linear Units (Shazeer2020Glu), dynamic unpadding (Zeng2022Boosting), and low-precision normalization. This engineering effort demonstrated that BERT pretraining can run significantly faster and more cost-effectively, while maintaining competitive downstream performance. In parallel, task-oriented pretraining strategies were explored, exemplified by RetroMAE (Xiao2022RetroMAE), which redesigned masked autoencoding to better support dense retrieval tasks using a BERT-style encoder. FlexiBERT (Tuli2023FlexiBERT) used neural architecture search to identify performant encoder configurations under different computational constraints, aiming to investigate architectural flexibility and efficiency trade-offs. Renewed interest in encoder-only models has led to modernized successors to BERT. ModernBERT (Warner2024ModernBERT) revisited the original BERT model, extending context length and updating architectural elements and design choices. NeoBERT (Breton2025NeoBERT) and OptiBERT (Dervishi2025TrainingOptiBERT) emphasized compute-optimal scaling and practical deployment. Together, these studies show a trend toward improving the efficiency and practicality of BERT-style encoder models, while maintaining strong performance on language understanding tasks.
@@ 2.2 Monolingual Turkish BERT Models
The success of transformer-encoder models motivated a series of BERT models known as BERTurk to be trained on Turkish data (Schweter2020BERTurk). Several variations of BERT, such as DistilBERT (Sanh2019DistilBERT), ConvBERTurk (Schweter2020BERTurkV2), and electra (Clark2020ELECTRA) were also trained on Turkish data by the same authors (Schweter2020BERTurk). These models were generally trained on a 3535GB corpus consisting of 4.44.4B tokens drawn from the Turkish oscar corpus (Abadji2022OSCAR), a Wikipedia dump [^1] An unspecified Wikipedia dump, presumed to date from approximately 2020., and various OPUS corpora (Tiedemann2012OPUS). Some models, like ConvBERTurk and electra, were also trained on the Turkish portion of the mC4 corpus (xue2021mt5massivelymultilingualpretrained; Schweter2020BERTurk), which is a cleaned version of the public web crawl data of Common Crawl. These models have been evaluated on various downstream tasks (such as part-of-speech tagging, named entity recognition, and question answering) where they generally outperform their multilingual counterparts (Schweter2020BERTurk; wu2019betobentzbecassurprising; Conneau2020XLM-R). Despite their popularity, these models were limited in transparency as they were not products of academic publications.
Following, two Turkish BERT models were introduced with academic publications. ytu-cosmos-BERT (Kesgin2023YTU) utilized an uncased Turkish tokenizer and was trained on a corpus of an unknown number of tokens reported to be around 7575GB in size, that consists of text collected from the web, Turkish Wikipedia and novels. TurkishBERTweet (Najafi2023TurkishBERTweet) was tailored for social media. It also employed an uncased tokenizer, specifically designed for social media texts, by the addition of special tokens that can encode special information in tweets like hashtags, mentions, and URLs. It was pretrained with a corpus of around 110110GB in uncompressed form, containing about 900 million tweets of 10 tokens on average, which totals to a sum of 99B tokens when extrapolated. Recently, a RoBERTa-based model for Turkish was published (ScheibleSchmitt2025SindBERT), which has been an improvement over the previous RoBERTa-based models for Turkish (TasRoberTURK2024) as it was trained on a larger corpus, using a larger batch size. However, it did not achieve a significant improvement over the previous models.